# Machine Unlearning: A Comprehensive Survey

### v1

@@ 4. Techniques of Centralized Unlearning
SISA (Bourtoule et al., 2021) Exact unlearning Samples Split learning Data and Model Partition 2021
GraphEraser (Chen et al., 2021a) Exact unlearning Graph nodes Split learning Data Partition 2021
Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split learning Partially retraining 2021
RecEraser (Chen et al., 2022) Exact unlearning Samples Split learning Balanced Data Partition 2022
ARCANE (Yan et al., 2022) Exact unlearning Samples Split learning Parition by Class 2022
HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split learning Tree ensemble learning 2021
DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split learning L-BFGS (Byrd et al., 1994) 2020
@@
(Ginart et al., 2019) Approximate unlearning Samples Certified Data Removal Quantized and DC-k-means 2019
Graph unl. (Chien et al., 2022) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian-based Unlearning VBI 2020
MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian-based Unlearning Monte cario-based 2022
BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian-based Unlearning MCMC 2022
@@
  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation (Chaudhuri et al., 2011), TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
@@
The naive retraining method is the most effective way to realize machine unlearning. However, it is expensive in both computation and storage. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. It is usually realized by unlearning based on the erased data only. In this section, we will introduce the popular techniques used in existing unlearning methods, in both exact unlearning and approximate unlearning. The dominant studies are summarized in Table 1, where the primary technique used in exact unlearning is split learning, and two primary techniques used in approximate unlearning are certified data removal and bayesian-based unlearning, respectively. The introduction order of these techniques will mainly follow the order of the unlearning techniques in Fig. 1.
4.1. Split Unlearning
Since most of the exact unlearning methods tried to split the full dataset into multi sub-sets and divide the model training process, we call this kind of unlearning technique split unlearning. The main process of split unlearning is illustrated in Fig. 5 (b). By contrast, we show the process of the naive retraining method in Fig. 5 (a). From Fig. 5 (a), we can see that when we unlearn the erased data by naive unlearning, we need two steps to realize it: first, remove the data from the original training dataset; second, retrain a new model based on the remaining dataset. Since it needs to store the whole training dataset and retrain from the remaining dataset, it often entails significant computational and stored overhead. Researchers saw the huge computation cost of naive retraining. They proposed many exact unlearning methods, and we summary as split unlearning because they were all based on data or model split techniques. As shown in Fig. 5 (b), the split unlearning technique can be summarized into four steps. Different from naive unlearning trained based on the whole remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards. All the constituent models are trained based on each split data slice. Then, in the second phase, when the unlearning request comes, they only need to remove the requested data from the split slice and retrain this slice’s constituent model in the third phase. In the last phase, the split unlearning aggregates the retrained and other constituent models together as a new unlearned model.
The first split unlearning is proposed by Cao and Yang (Cao and Yang, 2015). They split the original learning algorithms into a summation form. In a regular machine learning form, the model directly learns from the training dataset. However, in the summation form, they first train a small number of constituent models, which learn from several parts of the full trainset and then aggregate these intermediate models as the final learning model. So that when unlearning, they only need to retrain the constituent model that contains the information of erased data. It can efficiently speed up retraining time and reduce computation costs. In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many machine learning algorithms could be implemented in a summation form to reduce the retraining cost. The statistical query (SQ) learning (Kearns, 1998) guarantees the summation form. Although algorithms in Probably Approximately Correct (PAC) setting can be converted to the SQ learning setting, a huge amount of complex models, such as DNNs, cannot be converted to SQ learning efficiently.
@@ 4.1. Split Unlearning
Figure 5. (a) Naive unlearning. Only two steps: remove the erased data from the whole dataset and retrain the model based on the remaining dataset. (b) Split unlearning. It contains four steps: 1. split the original dataset into $n$ shards, 2. remove the erased data from the corresponding shard, 3. retrain the sub-model of this shard, 4. ensemble all sub-models as the final model.
@@
Then, Bourtoule et al. (Bourtoule et al., 2021) and Yan et al. (Yan et al., 2022) proposed advantaged methods unlearn samples suitable on deep neural networks. The primary idea of (Bourtoule et al., 2021; Yan et al., 2022) is also similar to the process shown in Fig. 5 (b). In (Bourtoule et al., 2021), Bourtoule et al. named their unlearning method the SISA training approach. SISA was implemented on deep neural networks, including two Conv. layers followed by two FC layers, Wide ResNet-1-1 and ResNet-50 models on multiple datasets, including MNIST, Purchase, SVHN, CIFAR100, Imagenet, and Mini-Imagenet. Unlike the original split unlearning dividing the dataset and transforming learning algorithms to summation form, Yan et al. proposed ARCANE (Yan et al., 2022), which transforms conventional ML into ensembling multiple one-class classification tasks. When many unlearning requests come, it can reduce retraining costs, which is not considered by previous work.
@@
Chen et al. extended exact unlearning methods to recommendation tasks and proposed RecEraser (Chen et al., 2022), which has similar architecture as split unlearning in Fig. 5 (b). RecEraser is tailored to recommendation systems, which can efficiently implement unlearning. Specifically, they designed three data partition schemes to split recommendation data into balanced pieces and propose an adaptive aggregation method based on an attention mechanism. They conducted the experimental evaluation on three real-world datasets, YELP2018, Movielens-1m, and Movielens-10m. These three datasets are widely used to assess the effectiveness and efficiency of recommendation models.
@@
Schelter et al. proposed HedgeCut (Schelter et al., 2021), which implemented machine unlearning on tree-based ML models in a split unlearning similar form. Tree-based learning algorithms are trained by recursively partitioning the training data, locally optimizing a criterion such as Gini gain (Wehenkel and Pavella, 1991). HedgeCut focuses on implementing fast retraining for these methods. Furthermore, they evaluated their method on five publicly available datasets on both accuracy and running time.
@@
Another method that is similar to split unlearning is Amnesiac Unlearning (Graves et al., 2021). The intuitive idea of Amnesiac Unlearning is to store the parameters of training batches and then subtract them when unlearning requests appear. In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{initial}$, where $E$ is the training epochs, and $B$ is the data batches. During the training process, they kept a list of $SB$, which records the batches containing the private data. The list could be in the form of an index of batches for each training example, an index of batches for each training category or any other information expected. When the unlearning request comes, a model using Amnesiac unlearning needs only to remove the updates from each batch $sb\in SB$ from the learned model $\theta_{M}$. As Graves et al. mentioned, using Amnesiac unlearning effectively and efficiently removes traces of the erased data that could be inferred via state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
4.2. Certified Data Removal
@@
Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition. Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction. After estimating the contribution of the erased data samples, they unlearn by subtracting these contributions with an updating bound from the unlearning model.
@@
In (Guo et al., 2020), Guo et al. proposed a certified data removal mechanism, which assumes removing the last training sample, $(x_{n},y_{n})$. Specifically, they defined a removal mechanism that approximately minimizes $\mathcal{L}(\theta;D^{\prime})$ with $D^{\prime}=D\backslash(x_{n},y_{n})$. The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\theta^{*}+\nabla l((\theta^{*})^{T}x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta^{*}$ by $H_{\theta^{*}}=\nabla^{2}L(\theta^{*};D^{\prime})$. Then they applied a one-step Newton update to the gradient impact of the removed point $(x_{n},y_{n})$ on the model $\theta$. Under their observation, they found that directly removing the Hessian contribution from the direction of the gradient will leak the privacy of the training data that was expected to be removed. They used the loss perturbation technique (Chaudhuri et al., 2011) to hide this information. It used a random linear term to perturb the empirical risk and ensures that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$. In(Mehta et al., 2022) and (Thudi et al., 2022), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020). Mehta et al. unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022). Thudi et al. (Thudi et al., 2022) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
@@
Another similar unlearning method is PUMA. In (Wu et al., 2022a), Wu et al. proposed a new data removal method through gradient re-weighting called PUMA, which also used the Hessian Vector Product (HVP) term. They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than a dot product between the pre-cached HVP term and individual gradient. When the unlearning request comes, they subtract the estimation of the erased samples to update the model parameters.
@@
Ginart et al. (Ginart et al., 2019) extended certified data removal to k-means clustering algorithms. They formulated the unlearning problem of efficiently removing personal data information from trained clustering models and proposed two different efficient deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means. In their work, both algorithms have theoretical guarantees and strong empirical results.
@@
DeltaGrad (Wu et al., 2020), an algorithm for fast retraining of SGD-based models, was proposed by Wu et al. to small changes of data, based on the idea of ”differentiating the optimization path” concerning the data and Quasi-Newton methods. They theoretically proved that their algorithm could approximate the right optimization path rapidly for the strongly convex objective. DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients exactly. After that, it only calculates the complete gradients for every first iteration. For other iterating rounds, it operates the L-BGFS algorithm (Byrd et al., 1994) to compute Quasi-Hessians approximating the true Hessians, keeping a set of updates at some prior iterations.
@@
Sekhari et al. (Sekhari et al., 2021) further given a strict separation between $\epsilon$-indistinguishable unlearning and differential privacy. Different from (Guo et al., 2020), in order to utilise tools of differential privacy (DP) for ML, the most straightforward manner is to ignore the special set of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that only relies on the learned algorithm $\mathcal{A}(D)$. In particular, the unlearning algorithm is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Notice the difference between (Guo et al., 2020) and (Sekhari et al., 2021). In the definition of (Guo et al., 2020), their $\epsilon$-indistinguishable unlearning is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{A}(D\backslash D_{e})$, but here is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighbouring datasets mean that for two datasets with an edit distance of $m$ samples. The guarantee of DP unlearning is stronger than the model distribution undistinguishable unlearning guarantee in (Guo et al., 2020), and thus it suffices to satisfy it. Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if it is private for datasets with edit distance $m$. Therefore, they derivate the bound on deletion capacity from the standard performance guarantees for DP learning. Furthermore, they determine that the existing unlearning algorithms can delete up to $\frac{n}{d^{1/4}}$ samples meanwhile still maintaining the performance guarantee w.r.t. the test loss, where $n$ is the size of the original trainset and $d$ is the dimension of trainset inputs.
@@
The above methods are all answering the problem of unlearning. Neel et al. (Neel et al., 2021) extended the unlearning definition to the update, which includes both ’add’, $D\cup\{z\}$, and ’delete’ as $D\backslash\{z\}$, where $z\in\mathcal{Z}$ is a data point. They follow the definition in (Guo et al., 2020) and define the similar $(\epsilon,\delta)$-indistinguishability. Moreover, they extend $(\epsilon,\delta)$-indistinguishability for both ’add’ and ’delete’ updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
4.3. Bayesian-based Unlearning

### v2

@@ 4.2. Detailed Techniques of Centralized Unlearning
SISA (Bourtoule et al., 2021) Exact unlearning Samples Split Unlearning Data and Model Partition 2021
Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split Unlearning Partially retraining 2021
GraphEraser (Chen et al., 2022b) Exact unlearning Graph nodes Split Unlearning Data Partition 2022
RecEraser (Chen et al., 2022a) Exact unlearning Samples Split Unlearning Balanced Data Partition 2022
ARCANE (Yan et al., 2022) Exact unlearning Samples Split Unlearning Parition by Class 2022
HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split Unlearning Tree ensemble learning 2021
DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split Unlearning L-BFGS (Byrd et al., 1994) 2020
ERASER (Hu et al., 2024a) Exact unlearning Samples Split Unlearning Inference Serving-Aware 2024
@@
(Warnecke et al., 2023) Approximate unlearning Samples Certified Data Removal Influence Theory 2023
Graph unl. (Chien et al., 2022a) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
Gif (Wu et al., 2023b) Approximate unlearning Graph nodes Certified Data Removal Influence function 2023
SUMMIT (Zhang et al., 2024) Approximate unlearning Graph nodes Certified Data Removal Multi-Objective Optimization 2024
EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian Unlearning VBI 2020
MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian Unlearning Monte cario-based 2022
BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian Unlearning MCMC 2022
@@
  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation, TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
@@
This section presents the popular techniques used in existing unlearning methods, in both exact and approximate unlearning studies. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. The dominant studies are summarized in Table 2, where the primary technique used in exact unlearning is split learning. Two primary techniques used in approximate unlearning are certified data removal and Bayesian-based unlearning. The introduction of these techniques will primarily follow the sequence of the unlearning techniques presented in Fig. 1.
4.2.1. Split Unlearning
@@
Figure 5. (a) Naive unlearning. There are only two steps: delete the specified samples from the whole dataset and retrain a model based on the remaining dataset. (b) Split unlearning. It contains four steps: 1. split the original dataset into $n$ shards, 2. remove the erased data from the corresponding shard, 3. retrain the sub-model of this shard, 4. ensemble all sub-models as the final model.
Since most exact unlearning methods attempted to partition the training dataset into multiple subsets and divide the ML model learning process, we call this kind of unlearning technique split unlearning. The main procedure of split unlearning is illustrated in Fig. 5 (b). By contrast, the process of the naive retraining method is shown in Fig. 5 (a), where we need two steps to realize it: first, delete the samples from the training dataset; second, retrain a new model using the remaining dataset. Since it needs to store the whole training dataset and retrain from the remaining dataset, it often entails significant computational and stored overhead. They proposed many exact unlearning methods to reduce the huge computation cost of naive retraining, and the majority of them are based on split learning techniques, either on data or model. As shown in Fig. 5 (b), the split unlearning technique can be summarized into four steps. Unlike naive unlearning trained based on the remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards. All the constituent models are trained based on each split data slice. Then, in the second phase, when the unlearning request comes, they only need to erase the requested samples from the split slice and retrain this slice’s constituent model in the third phase. In the last phase, the split unlearning aggregates the retrained and other constituent models together as a new unlearned model.
The first split unlearning is proposed by Cao and Yang (Cao and Yang, 2015). They split the original learning algorithms into a summation form. In a regular machine learning form, the model directly learns from the training dataset. However, in the summation form, they first train a small number of constituent models, which learn from several parts of the full trainset and then aggregate these intermediate models as the final learning model. So that when unlearning, they only need to retrain the constituent model that contains the information of erased data. It can efficiently speed up retraining time and reduce computation costs. In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many ML algorithms could be implemented in a summation form to reduce the retraining cost. The statistical query (SQ) learning (Kearns, 1998) guarantees the summation form. Although algorithms in the Probably Approximately Correct (PAC) setting can transform to the SQ learning setting, many complex models, such as DNNs, cannot be efficiently converted to SQ learning.
@@
Then, Bourtoule et al. (Bourtoule et al., 2021) and Yan et al. (Yan et al., 2022) proposed advantaged methods unlearn samples suitable on deep neural networks. The primary idea of (Bourtoule et al., 2021; Yan et al., 2022) is also similar to the process shown in Fig. 5 (b). In (Bourtoule et al., 2021), Bourtoule et al. named their unlearning method the SISA training approach. SISA can be implemented on deep neural networks, training multiple sub-neural networks based on divided sub-datasets. When the unlearning request comes, SISA retrains the model of the shard, which contains the information about the erased samples. SISA is effective and efficient as it aggregates all sub-models final prediction results rather than aggregates all these models. Unlike the original split unlearning dividing the dataset and transforming learning algorithms to summation form, Yan et al. proposed ARCANE (Yan et al., 2022), which transforms conventional ML into ensembling multiple one-class classification tasks. When many unlearning requests come, it can reduce retraining costs, which was not considered in previous work.
@@
Chen et al. (Chen et al., 2022a) extended exact unlearning methods to recommendation tasks and proposed RecEraser, which has similar architecture as split unlearning in Fig. 5 (b). RecEraser is tailored to recommendation systems, which can efficiently implement unlearning. Specifically, they designed three data division schemes to partition recommendation data into balanced pieces and created an adaptive aggregation algorithm utilizing an attention mechanism. They conducted the experiments on representative real-world datasets, which are usually employed to assess the effectiveness and efficiency of recommendation models.
@@
Besides the above popular ML models, Schelter et al. proposed HedgeCut (Schelter et al., 2021), which implemented machine unlearning on tree-based ML models in a split unlearning similar form. Tree-based learning algorithms are developed by recursively partitioning the training dataset, locally optimizing a metric such as Gini gain (Wehenkel and Pavella, 1991). HedgeCut focuses on implementing fast retraining for these methods. Furthermore, they evaluated their method on five publicly available datasets on both accuracy and running time.
@@
Another method that is similar to split unlearning is Amnesiac Unlearning (Graves et al., 2021). The intuitive idea of Amnesiac Unlearning is to store the parameters of training batches and then subtract them when unlearning requests appear. In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{\text{initial}}$, where $E$ is the training epochs, and $B$ is the data batches. In the model training process, they kept a list called $SB$, which records the batches holding the private data. This list could be formed as an index of batches for each training example, an index of batches for each category or any other information expected. When the unlearning request comes, a model using Amnesiac unlearning needs only to remove the updates from each batch $sb\in SB$ from the learned model $\theta_{M}$. As Graves et al. (Graves et al., 2021) stated, using Amnesiac unlearning effectively and efficiently removes the contribution of the erased samples that could be detected through state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
4.2.2. Certified Data Removal
Figure 6. The approximate unlearning by certified data removal. The unlearning algorithm $\mathcal{U}$ includes estimating the influence of specified data and removing the estimation from trained models. The unlearned model is expected to approach the retrained model.
@@ 4.2.2. Certified Data Removal
Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition (Dwork et al., 2014). An example is presented in Figure 6. Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction. After estimating the impact of the erased data samples, they unlearn by subtracting these impacts with an updating bound from the unlearning model.
@@
In (Guo et al., 2020), Guo et al. proposed a certified data removal method, which assumes removing the last training sample, $(x_{n},y_{n})$. Specifically, they defined a removal mechanism that approximately minimizes $\mathcal{L}(\theta;D^{\prime})$ with $D^{\prime}=D\backslash(x_{n},y_{n})$. The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\nabla\ell(\theta^{T}\cdot x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta$ by $H_{\theta}=\nabla^{2}L(\theta;D^{\prime})$. Then, they applied a one-step Newton update to the model parameters impact of the erased point $(x_{n},y_{n})$ on the model $\theta$. Under their observation, they found that directly removing the Hessian contribution from the gradient will reveal the private information of the erased data. They used the loss perturbation technique (Chaudhuri et al., 2011) to hide this information. It used a random linear term to perturb the empirical risk and ensure that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$. In(Mehta et al., 2022) and (Thudi et al., 2022a), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020). Mehta et al. (Mehta et al., 2022) unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022). Thudi et al. (Thudi et al., 2022a) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
@@
Another similar unlearning method is PUMA. In (Wu et al., 2022a), Wu et al. proposed a new data removal method through gradient re-weighting called PUMA, which also used the Hessian Vector Product (HVP) term. They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than one dot product between the pre-cached HVP term and individual gradient. When the unlearning request comes, they subtract the estimate of the erased samples to revise the model.
@@
Ginart et al. (Ginart et al., 2019) extended certified data removal to k-means clustering algorithms. They formulated the unlearning problem of efficiently removing personal data information from trained clustering models. They offered two different deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means. In their work, both algorithms have theoretical guarantees and strong empirical results.
@@
To retrain SGD-based models fast, DeltaGrad was proposed by Wu et al. (Wu et al., 2020) to unlearn small changes of data inspired by the idea of ”differentiating the optimization path” concerning the training dataset and Quasi-Newton methods. They theoretically proved that their algorithm could approximate the right optimization path rapidly for the strongly convex objective. DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients precisely. After that, it only calculates the complete gradients for every first iteration. For other iterating rounds, it operates the L-BGFS algorithm (Byrd et al., 1994) to compute Quasi-Hessians approximating the true Hessians, keeping a set of updates at some prior iterations.
@@
For a deeper understanding of certified machine unlearning, Sekhari et al. (Sekhari et al., 2021) further given a strict separation between $\epsilon$-indistinguishable unlearning and differential privacy. Different from (Guo et al., 2020), in order to utilize tools of differential privacy (DP) for ML, the most straightforward manner is to forget the special dataset of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that solely relies on the learned algorithm $\mathcal{A}(D)$. In particular, the unlearning method is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Notice the difference between (Guo et al., 2020) and (Sekhari et al., 2021). In the definition of (Guo et al., 2020), their $\epsilon$-indistinguishable unlearning is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{A}(D\backslash D_{e})$, but here is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighboring datasets mean that for two datasets with an edit distance of $m$ samples. The guarantee of DP unlearning is more powerful than the model distribution undistinguishable unlearning in (Guo et al., 2020), and therefore, it suffices to satisfy it. Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if they are private for datasets with the distance $m$. Therefore, they derivate the bound on deletion capacity from the standard performance guarantees for DP learning. Furthermore, they determine that the existing unlearning algorithms can delete up to $\frac{n}{d^{1/4}}$ samples meanwhile still maintaining the performance guarantee w.r.t. the test loss, where $n$ is the size of the original trainset, and $d$ is the dimension of trainset inputs.
@@
The aforementioned methods are trying to address the basic unlearning problem from a certified data removal perspective. Neel et al. (Neel et al., 2021) extended the definition of unlearning to include updates, which encompass both ”add” ($D\cup\{z\}$) and ”delete” ($D\backslash\{z\}$), where $z=(x,y)$ and $z\in\mathcal{Z}$ is a data point. They follow the definition in (Guo et al., 2020) and define similar $(\epsilon,\delta)$-indistinguishability. Furthermore, they extend $(\epsilon,\delta)$-indistinguishability to both ”add” and ”delete” updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
4.2.3. Bayesian-based Unlearning